Cariboo Manufacturing Company incurred a joint cost of $759,000 in the production of X and Y in a joint process. Presently, 2,500 of X and 2,100 of Y are being produced each month. Management plans to decrease X's production by 650 units in order to increase the production of Y by 920 units. Additionally, this change will require minor modifications, which will add $51,520 to the joint cost. This cost is entirely attributable to product Y. What is the amount of the joint costs allocable to X and Y before changes to existing production assuming Cariboo allocates their joint costs according to the proportion of X and Y produced?

Product X Product Y
A. $ 346,500 $ 412,500
B. $ 339,706 $ 419,924
C. $ 348,529 $ 410,471
D. $ 412,500 $ 346,500
Answer #1

Total Unit produced before changes = 2500+2100 = 4600

Join Cost allocation = Join Cost/Total Unit produced*Unit produced product X

Product X = 759000*2500/4600 = 412500

Product Y = 759000*2100/4600 = 346500

So answer is d)